Indicate the equation of the given line in standard form.
The line with slope 9/7 and containing the midpoint of the segment whose endpoints are (2,-3) and (-6,5)

Answer by stanbon(75887)

Indicate the equation of the given line in standard form.
The line with slope 9/7 and containing the midpoint of the segment whose endpoints are (2,-3) and (-6,5)
-----
The midpoint is [(2-6)/2 , (5-3)/2] = (-2,1)
-----
Form: y = mx+b
1 = (9/7)(-2) + b
1 = -18/7 + b
b = 25/7
---------------
Equation:
y = (9/7)x + 25/7
=====================
Cheers,
Stan H.
=====================